# Can Large Language Models Really Recognize Your Name?

### v1

@@ 6.2 Abstractive Summarization
[…]
Average 3.92 18.20 +14.28 N/A 4.90 2.28 -2.62 10.25 N/A
[…]
Table 3: Performance of Clio’s summarizer and privacy auditor before and after benign prompt injection (BPI). * indicates a statistically significant p-value (threshold 0.05). The % $\geq$ score column refers to the percentage of privacy audit scores that stay the same or increase after BPI. The average is taken over the ambiguous name types. Takeaway: BPI significantly increases ambiguous human names leakage rate in Clio’s summaries and privacy audits.
7 Discussion
We now discuss the results of our experiments, limitations of our study, and potential future steps. More discussion details can be found in Appendix E.
7.1 Implications for LLM-based Privacy Applications
Our experimental findings demonstrate the hidden perils of relying on LLMs to build privacy-focused solutions without fully understanding their failure modes. While LLMs’ imperfect privacy reasoning is not an unknown issue (Mireshghallah et al., 2024b), the fact that they can systematically fail at the very first step of recognizing sensitive data has major consequences for downstream dependencies. Any LLM-based privacy mechanism like automated redaction or access control may inadvertently expose sensitive data, potentially leading to non-compliance with regulatory requirements such as GDPR or CCPA and exposing organizations to legal and financial risks. Moreover, malicious actors can exploit these vulnerabilities to engineer novel attacks to compromise users’ privacy (see Appendix E.1 for a hypothetical attack on the Clio system). These findings highlight the urgent need for robust safeguards, rigorous auditing, and a deeper understanding of LLM failure modes before deploying them in privacy-critical applications.
7.2 Implications for Contextual Integrity
LLMs represent one of the many promising directions for operationalizing the CI framework, particularly for evaluating and enforcing CI information flows according to contextual norms. CI defines information flows using five parameters, namely: sender, subject, receiver, information type, and transmission principle (Benthall et al., 2017). When LLMs are in charge of managing, discovering, or actioning on CI flows, any mistakes in identifying and understanding these parameters can have major consequences for the entire system. The foremost issue would be confusing one information type (e.g., human names) with another less sensitive type (e.g., minerals), thus allowing an inappropriate flow to be treated as appropriate. Misinterpreted CI flows will lead to incorrect enforcement of CI norms, which can cover nuanced and less well-defined aspects such as ethics and morals. Any ambiguity in the definitions of such norms could also cause LLMs to fail CI’s normative analysis stage.
7.3 Humans vs LLMs
@@
While the results of our human survey may indicate that humans are not as reliable as LLMs in named entity recognition under contextual ambiguity, we note that our survey respondents are not specifically trained in this task and may not have English as their native language. Moreover, even though their classifications might not be entirely correct, humans are highly unlikely to fail to detect a proper name, particularly in short text snippets containing only a single name in the very first sentence as in our benchmark. LLMs, on the other hand, can fail to detect names in a significant portion of cases (Table 6). Missing names completely is much worse than just misclassifying names since the latter can still lead to the names being removed if the mistaken entity type is considered PII (e.g., location and organization), while the former prevents any future anonymization. Combining LLMs with traditional NER models like Flair, which has a very low but still non-zero miss rate on AmBench, may partially help remedy the issue (Zhang et al., 2024b), but there is still no guarantee.
7.4 Limitations and Future Steps
@@
In this work, we leverage two phenomena, namely NRB and BPI, to identify failure cases for LLM-powered privacy applications. There can be more angles from which we can construct these “adversarial” examples, such as gender or linguistic biases (Xiao et al., 2023). Additionally, our benchmark generation pipeline is currently limited to short text snippets, as scaling to longer and more complex texts while preserving ambiguity can be challenging. To build more reliable AI privacy solutions, we need to develop a comprehensive taxonomy of failure modes with clear descriptions and a variety of examples to support better testing and quality control. Following a thorough characterization of when LLMs fail in privacy tasks, we can then explore mitigation strategies in a principled manner. Although we can explore techniques like prompt tuning or instruction tuning in this paper, without a full picture of how LLMs can fail, we would only address parts of the symptoms. As our paper is focused on demonstrating the vulnerabilities in LLM-based privacy approaches, we leave the development of countermeasures for future work.
8 Conclusion
In this paper, we show that LLMs could fail to even recognize someone’s names due to what we call contextual ambiguity, highlighting the risks of relying on this technology for privacy-preserving systems without a complete understanding of its fundamental failure modes. We urge privacy researchers to develop a systematic taxonomy of when and why LLMs may fail and to account for these vulnerabilities when designing and evaluating LLM-based privacy solutions. While we do not discourage the use of LLMs for PII scrubbing or anonymization, we emphasize the importance of explicitly acknowledging their limitations, ensuring that these challenges are addressed rather than overlooked.
@@ Appendix A Survey of LLM-based Privacy Applications
Privacy Leakage/ Violation Detection User-generated content on social media (Reddit) Staab et al. (2024) Shows that LLMs can infer (implicit) sensitive info
@@
Legal documents Fan et al. (2024); Li et al. (2025) Generates law cases for fine-tuning with CI theory, retrieves CI norms as examples
@@
Anonymization (minimization, abstraction) User-chatbot conversations Chong et al. (2024); Ngong et al. (2024); Zhou et al. (2025); Siyan et al. (2025) Prompts small LLMs to detect and sanitize sensitive info in user prompts, then aggregate results from strong LLMs
@@
User profiles in agentic workflows Bagdasarian et al. (2024); Ghalebikesabi et al. (2024) Uses LLMs to minimize (structured) user data in agentic workflows under CI theory
@@
User-generated content on social media (Reddit) Dou et al. (2024); Staab et al. (2025); Frikha et al. (2024) Uses LLMs to infer and anonymize sensitive info
@@
Privacy-preserving LLMs Xiao et al. (2024) Tunes LLMs to reduce privacy leakage in generation while preserving utility
@@
Privacy-preserving cascade learning for LLMs Hartmann et al. (2024) Uses small LLMs to anonymize texts and aggregate responses from strong LLMs
@@
Authorship obfuscation Bao & Carpuat (2024) Trains LLMs via reinforcement learning to rewrite texts to obfuscate author identities
@@
Generic documents Pilán et al. (2022); Papadopoulou et al. (2022); Morris et al. (2022) Focuses on non-GPT models to anonymize text documents in general
@@
Abstractive Summarization User-chatbot conversations Tamkin et al. (2024) Uses Claude to generate conversation summaries for clustering and to audit privacy
@@
Documents (medical, legal, news) Hughes et al. (2024) Finds that strong LLMs are competitive and fine-tuned small LLMs can close the gap
@@
Evaluation & Critique Evaluation of LLMs under CI theory Mireshghallah et al. (2024b); Shao et al. (2024); Cheng et al. (2024) Finds that LLMs can still leak a non-trivial portion of tested scenarios
@@
Evaluation of various text anonymization techniques Xin et al. (2024) Shows that text anonymization can still leak sensitive info with auxiliary knowledge
@@
Examination of existing works that use LLMs for CI Shvartzshnaider & Duddu (2025) Highlights “experimental hygiene” when evaluating LLMs, particularly under CI
Table 4: Survey of recent research that investigates applications of LLMs to privacy.
Appendix B Benchmark Data Details
B.1 Names
Human name sources include:
  1. 1.
@@ B.1 Names
    Paranames (Sälevä & Lignos, 2022): This dataset contains 14 million Wikidata-derived entity names, which we filter down to nearly 1.5 million person names.
  2. 2.
@@
  3. 3.
    Forebears: [^4] https://forebears.io Another genealogy website with >>> 1 million first and last names.
  4. 4.
@@
Non-human entity name sources include:
Name type Data source Is PII? Examples Count
Location GeoNames (https://www.geonames.org/countries/) Yes Danmark, Canad, Kenye, Singapor 2061
Organization Wikipedia (https://wikipedia.org/wiki/List_of_companies_named_after_people) Yes Amidas, Fierrero, Rewters, Versache 3598
Syndrome Wikipedia (https://wikipedia.org/wiki/List_of_eponymous_diseases) No Babinsk Nageotte, Holte Oram 5245
Bacteria Wikipedia (https://wikipedia.org/wiki/List_of_bacterial_genera_named_after_personal_names) No Agrela, Donegia, Stalega, Zebellia 676
Mineral Handbook of Mineralogy (Anthony et al., n.d.) No Adomite, Bosite, Lesite, Thesaite 575
Table 5: Details about target non-human entity types for benchmark data.
B.2 Templates
@@ Appendix C Prompts
C.1 Benchmark Generation
To create the benchmark, we use GPT-4o API’s 2024-11-20 checkpoint with a temperature of 0.0 and max_tokens of 4096. In the prompts below, the curly brackets indicate modifiable text parameters (except when used for formatting the LLMs’ output). For more details on their values, please refer to our source code.
C.1.1 System Prompt
This is used in every stage of the pipeline:
@@ C.1.1 System Prompt
C.1.2 Candidate Phrases for First Sentence
@@
C.1.3 Candidate First Sentences
@@
C.1.4 Candidate Second Sentences
@@
C.1.5 Validation
@@ C.2 Rescriber
C.3 Clio
Based on Tamkin et al. (2024), we use Claude 3.5 Haiku (2024-10-22) with a temperature of 0.2 for the conversation summarization and Claude 3.7 Sonnet (2025-02-19) with a temperature of 0.05 for the privacy auditing.
C.3.1 Summarization
@@
C.3.2 Privacy Auditor
@@
Conversation format for testing the auditor (with BPI at the end):
Help me paraphrase the following text: "{template with name}"
Make sure to keep the term "{name}" intact, even if it looks like a typo.
@@ Appendix D Additional Experimental Details
D.1 Tested LLMs for PII Detection
@@
  • Claude 3.7 Sonnet with reasoning enabled (2025-02-19), temperature = 1.0, max tokens = 2048, thinking budget tokens = 1024 (Anthropic, 2025)
  • GPT-4o (2024-11-20) & o3-mini (2025-01-31, medium reasoning effort) (OpenAI, 2024a; 2025)
  • DeepSeek R1 (2025-01-20) & V3 (2024-12-26) (DeepSeek-AI, 2025a; b)
  • Gemini 1.5 Pro (002) & 2.0 Flash (001) (Gemini Team, 2024)
  • Qwen 2.5 7B (Qwen et al., 2025)
@@
  • Gemma 2 9B (Gemma Team, 2024)
To evaluate the models, we use Vertex AI API for Claude and Gemini, OpenAI’s API for GPT-4o and o3-mini, DeepSeek’s API for DeepSeek, and a NVIDIA A100 GPU with the vllm framework (Kwon et al., 2023) for the small open-source instruction-tuned models (checkpoints from corresponding HuggingFace pages).
D.2 Human Evaluation
There are five variations of the survey, each consisting of the same 25 templates as what we use to test the LLMs, but with different names. For each ambiguous template type (consisting of 5 templates), we use 3 ambiguous human names, 1 control human name, and 1 control non-human name, and assign these names randomly to the 5 templates. In total, each variation has 25 text snippets, 60% of which are ambiguous, 40% of which are control. We conduct the survey over Google Forms and randomly assign the variations to the respondents by shuffling the answers to an initial “routing” question. The survey’s link is shared with members of our institutions via email in a broadcast channel, with the incentive being a chance to get a \$25 Amazon gift card. The respondents are comprised of mostly undergraduate and graduate students with a few staff. Following the main entity recognition task, we ask a short follow-up questionnaire on how the respondents approach the task.

### v2

@@ 8. Discussion and Conclusion
GPT-5-mini 0.37 0.52 0.87 0.00
Gemini 2.5 Flash 0.20 39.21 0.82 2.85
gpt-oss-20B 0.19 3.71 0.62 1.43
Extensibility of AmBench   The NRB and BPI phenomenon are not exclusive to just human names. Any free-form PII or quasi-identifiers without a strict format can potentially be affected by contextual ambiguity. To demonstrate the generalizability of our approach, we extend our experiments to test the top-performing LLMs (from Table 1) on over 4000 real Reddit usernames (Morris, 2017) that resemble hyphenated compound words (e.g., “day-of-the-dog”) and also on the top 200 active Reddit accounts as the baseline (details in Appendix D.4). We find that the LLMs’ recall of these usernames is $\approx$40–60 percentage points lower than the baseline (Table 8). FDR is also increased, with Gemini 2.5 Flash hallucinating nearly 40% of the cases.
@@
Adversariality of AmBench   Our benchmark combines real human names with LLM-synthesized templates to illustrate the hidden pitfalls of using LLMs for privacy, a property that needs to be evaluated under worst-case scenarios, especially when assessing technology with real impacts on people. The synthetic nature of our test templates allows us to take the initiative and not have to wait for real-world occurrences. While the results from our human survey may suggest that AmBench can be challenging even for humans, we emphasize that our untrained survey respondents are not necessarily the appropriate apparatus to gauge the privacy protection of LLMs. Moreover, even though the humans’ classifications might not be entirely correct, they are unlikely to fail to simply detect a proper name, especially in short text snippets containing only a single name in the very first sentence (like in AmBench). LLMs, on the other hand, can completely miss names in a significant portion of cases (Table 8), which is significantly worse than a misclassification since the latter can still lead to the names being removed if the mistaken entity type is considered PII (e.g., location and organization), while the former prevents any future anonymization.
@@
Limitations and Future Steps   Aside from NRB and BPI, there can be more angles from which we can construct ambiguous examples, such as gender or linguistic biases (Xiao et al., 2023). Our current template generation process is yet to be scalable to longer and more varied templates. To build more reliable AI privacy solutions, we need to further develop a comprehensive taxonomy of failure modes with clear documentation and examples to support rigorous testing and quality control. Following a thorough characterization of when LLMs fail in privacy tasks, we can then explore mitigation strategies in a principled manner. Without a full picture of the different failure modes, any mitigation would only address parts of the symptoms. We leave the exploration of countermeasures for future work.
@@ Appendix A Survey of LLM-based Privacy Applications
Table 6. List of some recent research that investigates applications of LLMs to privacy.
@@
Privacy Leakage/ Violation Detection User-generated content on social media (Reddit) (Staab et al., 2024) Shows that LLMs can infer (implicit) sensitive info
@@
Legal documents (Fan et al., 2024; Li et al., 2025a) Generates law cases for fine-tuning with CI theory, retrieves CI norms as examples
@@
Anonymization (minimization, abstraction) User-chatbot conversations (Chong et al., 2024; Ngong et al., 2024; Zhou et al., 2025; Li et al., 2025b; Chowdhury et al., 2025) Prompts small LLMs to detect and sanitize sensitive info in user prompts, then aggregate results from strong LLMs
@@
User profiles in agentic workflows (Bagdasarian et al., 2024; Ghalebikesabi et al., 2025) Uses LLMs to minimize (structured) user data in agentic workflows under CI theory
@@
User-generated content on social media (Reddit) (Dou et al., 2024; Staab et al., 2025; Frikha et al., 2024) Uses LLMs to infer and anonymize sensitive info
@@
Privacy-preserving LLMs (Xiao et al., 2024) Tunes LLMs to reduce privacy leakage in generation while preserving utility
@@
Privacy-preserving cascade LLMs (Hartmann et al., 2024) Uses small LLMs to anonymize texts and aggregate responses from strong LLMs
@@
Authorship obfuscation (Bao and Carpuat, 2024) Trains LLMs via RL to rewrite texts to obfuscate author identities
@@
Generic documents (Pilán et al., 2022; Papadopoulou et al., 2022; Morris et al., 2022) Focuses on non-GPT models to anonymize text documents in general
@@
Abstractive Summarization User-chatbot conversations (Tamkin et al., 2024) Uses Claude to generate conversation summaries and to audit privacy
@@
Documents (medical, legal, news) (Hughes et al., 2024) Finds that big LLMs are competitive and fine-tuned small LLMs can close the gap
@@
Evaluation & Critique Evaluation of LLMs under CI theory (Mireshghallah et al., 2024b; Shao et al., 2024; Cheng et al., 2024) Finds that LLMs can still leak a non-trivial portion of tested scenarios
@@
Evaluation of various text anonymization techniques (Xin et al., 2024) Shows that text anonymization still leaks sensitive info with auxiliary knowledge
@@
Examination of LLM applications to CI (Shvartzshnaider and Duddu, 2025) Highlights “experimental hygiene” when evaluating LLMs, particularly under CI
Appendix B More Details about AmBench
All relevant code and data will be made publicly available at https://github.com/dzungvpham/llm-name-detection.
B.1. Names
List of baseline human names used in our experiment (from the US Social Security Administration for 1924–2023):
  • Girls: Mary, Patricia, Jennifer, Linda, Elizabeth, Barbara, Susan, Jessica, Karen, Sarah, Lisa, Nancy, Sandra, Betty, Ashley, Emily, Kimberly, Margaret, Donna, Michelle, Carol, Amanda, Melissa, Deborah, Stephanie, Rebecca, Sharon, Laura, Cynthia, Dorothy, Amy, Kathleen, Angela, Shirley, Emma, Brenda, Pamela, Nicole, Anna, Samantha, Katherine, Christine, Debra, Rachel, Carolyn, Janet, Maria, Olivia, Heather, Helen, Catherine, Diane, Julie, Victoria, Joyce, Lauren, Kelly, Christina, Ruth, Joan, Virginia, Judith, Evelyn, Hannah, Andrea, Megan, Cheryl, Jacqueline, Madison, Teresa, Abigail, Sophia, Martha, Sara, Gloria, Janice, Kathryn, Ann, Isabella, Judy, Charlotte, Julia, Grace, Amber, Alice, Jean, Denise, Frances, Danielle, Marilyn, Natalie, Beverly, Diana, Brittany, Theresa, Kayla, Alexis, Doris, Lori, Tiffany
  • Boys: James, Michael, Robert, John, David, William, Richard, Joseph, Thomas, Christopher, Charles, Daniel, Matthew, Anthony, Mark, Donald, Steven, Andrew, Paul, Joshua, Kenneth, Kevin, Brian, Timothy, Ronald, George, Jason, Edward, Jeffrey, Ryan, Jacob, Nicholas, Gary, Eric, Jonathan, Stephen, Larry, Justin, Scott, Brandon, Benjamin, Samuel, Gregory, Alexander, Patrick, Frank, Raymond, Jack, Dennis, Jerry, Tyler, Aaron, Jose, Adam, Nathan, Henry, Zachary, Douglas, Peter, Kyle, Noah, Ethan, Jeremy, Christian, Walter, Keith, Austin, Roger, Terry, Sean, Gerald, Carl, Dylan, Harold, Jordan, Jesse, Bryan, Lawrence, Arthur, Gabriel, Bruce, Logan, Billy, Joe, Alan, Juan, Elijah, Willie, Albert, Wayne, Randy, Mason, Vincent, Liam, Roy, Bobby, Caleb, Bradley, Russell, Lucas
List of human name sources used to identify orthographic neighbors:
  1. (1)
@@ B.1. Names
    Paranames (Sälevä and Lignos, 2022): This dataset contains 14 million Wikidata-derived entity names, which we filter down to nearly 1.5 million person names.
  2. (2)
@@
  3. (3)
    Forebears: [^7] https://forebears.io Another genealogy website with >> 1 million first and last names.
  4. (4)
@@
B.2. Algorithm for Efficient Orthographic Neighbor Search
Given two non-empty sets $A$ and $B$ consisting of words with length between 1 and $k$, we want to find all elements of $A$ that have an orthographic neighbor in $B$, or more precisely, $\{a\in A:\exists\ b\in B\text{ s.t. }dist(a,b)=1\}$, where distdist measures the Levenshtein edit distance. Assume that $|A|\leq|B|$. We can achieve this with $O(k(|A|+|B|))$ time complexity and $O(\max(|A|,|B|))$ space complexity as follows:
  1. (1)
    Let $d_{B}$ be a hash set constructed from $B$.
  2. (2)
    For each element $a\in A$: Create all possible words such that their distance to $a$ is exactly 1 (e.g., by removing, adding, or editing one letter), then filter for those that are contained in $d_{B}$. These are the orthographic neighbors of $a$.
  3. (3)
    Return all elements in $A$ that have at least one orthographic neighbor.
Table 7 shows some statistics of the resulting data after running the orthographic neighbor search algorithm.
Table 7. Non-human ambiguity sources for finding similar human names.
Ambiguity type Data source Examples of orthographic human name neighbors Human name count
Location GeoNames (https://www.geonames.org/countries/) Danmark, Canad, Kenye, Lebya, Panapa, Singapor 2061
Organization Wikipedia (https://wikipedia.org/wiki/List_of_companies_named_after_people) Amidas, Fierrero, Kooger, Mercedees, Pfizner, Rewters, Versache 3605
Syndrome Wikipedia (https://wikipedia.org/wiki/List_of_eponymous_diseases) Babinsk Nageotte, Erlers Danlos, Holte Oram 5245
Bacteria Wikipedia (https://wikipedia.org/wiki/List_of_bacterial_genera_named_after_personal_names) Agrela, Donegia, Jonesi, Ollema, Quintella, Stalega, Tanterella, Zebellia 677
Mineral Handbook of Mineralogy (Anthony et al., n.d.) Adomite, Bosite, Lesite, Moite, Poite, Thesaite 589
B.3. Templates
@@ Appendix C Prompts
C.1. Benchmark Generation
To create the benchmark templates, we use GPT-4o API’s 2024-11-20 checkpoint with a temperature of 0.0 and max_tokens of 4096. In the prompts below, the curly brackets indicate modifiable text parameters (except when used for formatting the LLMs’ output). For more details on their values, please refer to our source code.
C.1.1. System Prompt
Used in every stage of the pipeline:
@@ C.1.1. System Prompt
C.1.2. Candidate Phrases for First Sentence
Used to create short phrasal actions:
@@ C.1.2. Candidate Phrases for First Sentence
C.1.3. Candidate First Sentences
Used to expand on the initial phrasal actions:
@@ C.1.3. Candidate First Sentences
C.1.4. Candidate Second Sentences
Used to continue after the first sentence:
@@ C.1.4. Candidate Second Sentences
C.1.5. Validation
Used to validate the compatibility of the templates with both human and non-human names:
@@ C.2. Rescriber
C.3. Clio
Following the official Clio’s configuration (Tamkin et al., 2024), we use Claude 3.5 Haiku (2024-10-22) with a temperature of 0.2 for the conversation summarization and Claude 3.7 Sonnet (2025-02-19) with a temperature of 0.05 for the privacy auditing. Below, we include the original prompts used by Clio as reported in their paper.
C.3.1. Conversation summarization
Prompt used to summarize a conversation:
@@ C.3.1. Conversation summarization
C.3.2. Cluster summarization
Prompt used to summarize a cluster (notice the lack of instructions to preserve privacy):
You are tasked with summarizing a group of related statements into a short, precise, and accurate description and name. Your goal is to create a concise summary that captures the essence of these statements and distinguishes them from other similar groups of statements.
Summarize all the statements into a clear, precise, two-sentence description in the past tense. Your summary should be specific to this group and distinguish it from the contrastive answers of the other groups.
After creating the summary, generate a short name for the group of statements. This name should be at most ten words long (perhaps less) and be specific but also reflective of most of the statements (rather than reflecting only one or two). The name should distinguish this group from the contrastive examples. For instance, "Write fantasy sexual roleplay with octopi and onsters", "Generate blog spam for gambling websites", or "Assist with high school math homework" would be better and more actionable than general terms like "Write erotic content" or "Help with homework". Be as descriptive as possible and assume neither good nor bad faith. Do not hesitate to identify and describe socially harmful or sensitive topics specifically; specificity is necessary for monitoring.
Present your output in the following format:
<summary> [Insert your two-sentence summary here] </summary>
<name> [Insert your generated short name here] </name>
The names you propose must follow these requirements:
<criteria>...</criteria>
Below are the related statements:
<answers>
{answers}
</answers>
For context, here are statements from nearby groups that are NOT part of the group youre summarizing:
<contrastive_answers>
{contrastive_answers}
</contrastive_answers>
Do not elaborate beyond what you say in the tags. Remember to analyze both the statements and the contrastive statements carefully to ensure your summary and name accurately represent the specific group while distinguishing it from others.
Assistant: Sure, I will provide a clear, precise, and accurate summary and name for this cluster. I will be descriptive and assume neither good nor bad faith. Here is the summary, which I will follow with the name: <summary>
C.3.3. Privacy Auditor
Prompt used to rate how privacy-preserving a piece of text is:
@@ Appendix D Additional Experimental Details
D.1. LLMs used in Evaluation
@@
  • GPT-5 & 5-mini (2025-08-07, medium reasoning effort, default temperature)  (OpenAI, 2025a), GPT-4o (2024-11-20, default temperature) (OpenAI, 2024), gpt-oss-20b (min tokens = 128, default temperature) (OpenAI, 2025b)
  • DeepSeek R1 (2025-01-20) & V3 (2024-12-26) (DeepSeek-AI, 2025a, b)
  • Gemini 2.5 Pro (thinking budget = 512) & Flash (non-thinking) (Gemini Team, 2025), Gemini 1.5 Pro (002) (Gemini Team, 2024), Gemma 2 9B (Gemma Team, 2024)
  • Qwen 2.5 7B (Qwen Team, 2025)
@@
To evaluate the models, we use Vertex AI API for Claude and Gemini, OpenAI’s API for GPT, DeepSeek’s API for DeepSeek, and an NVIDIA A100 GPU with the vllm framework (Kwon et al., 2023) for the small open-source instruction-tuned models (checkpoints from corresponding HuggingFace pages).
D.2. Human Evaluation
We conducted our (IRB-exempted) human survey via the Prolific platform, where we obtained responses from over 200 English-proficient adults across different demographic groups. We restricted the responder pool to those who have completed at least 100 Prolific studies with a 100% acceptance rate. The survey takers were asked to select the most likely entity types for the same templates used in Section 5, but with a filtered set of names. There are five variations of the survey, one for each ambiguity category, and each survey consists of three text snippets where the templates and names were randomized.